OpenStudy (anonymous):

Complex numbers: 3+i/3-i: complex numbers, please help!!

OpenStudy (anonymous):

multiply top and bottom by the conjugate of the bottom. http://www.mathwords.com/c/complex_conjugate.htm

OpenStudy (anonymous):

So, the answer is 8/10

OpenStudy (anonymous):

no... you should get \[\frac{ 3+i }{ 3-i }\cdot\frac{ 3+i }{ 3+i }=\frac{ 3\cdot3-1\cdot 1+ (3\cdot1 + 3\cdot 1)i}{ 3^2+1^2 }=\frac{ 8+6i }{ 10 }=\frac{ 4 }{ 5 }+\frac{ 3 }{ 5 }i\]